OpenStudy (anonymous):

find the coordinates of the midpoint of the segment whose endpoints are (-4 6) and (-8 -2) answer -6,2

OpenStudy (anonymous):

use formula of length \[\sqrt{a ^{2}+b ^{2}}\]

OpenStudy (evonhowell):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you. I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school. I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. If you have ANY questions about the site Please just let Me Ambassador @EvonHowell or any of the other Ambassadors know We Ambassadors Have the Purple A next to our name, And if there is a problem like someone cheating or being rude do Not hesitate to contact Me or any of the Ambassadors and we will do what we can to help the situation IF we can't get the situation under control We will call for a Moderator These are the people with the purple Square around their names. If you want us to get somewhere quickly just Tag us by putting the @ symbol then Type our names NO SPACES BUT DON'T do this to much! Have a fabulous Day!!!! WELCOME to the OpenStudy Community. Ambassador EvonHowell

OpenStudy (snuggielad):

oh my gosh...evan...delete that...that is way to much...it is overwelming to him... he is asking for help not a long explanation...if you overwelm him he will never come back..there is a time and a place for that...its not here or now

OpenStudy (anonymous):

find AB=B-A , where A and B are points given

OpenStudy (anonymous):

midpoint formula: \[\Large (x,y)=(\frac{x_{1}+x_{2}}{2}, \frac{y_{1}+y_{2}}{2})\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I am a girl not a boy .so she